what is the diameter of a 6ft tank that can hold 9 cubic ft
Answer by mananth(16949)

volume = 9 cu.ft
height = 6 ft
...
V= pi*r^2*h
9=pi*r^2*6
r^2=9/6*pi
r^2= 4.7
r=2.14
diameter = 4.28 ft
...
m.ananth@hotmail.ca